# When Avatars Have Personality: Effects on Engagement and Communication in Immersive Medical Training

### v1

@@ Abstract
While virtual reality (VR) excels at simulating physical environments, its effectiveness for training complex interpersonal skills is limited by a lack of psychologically plausible virtual humans. This is a critical gap in high-stakes domains like medical education, where communication is a core competency. This paper introduces a framework that integrates large language models (LLMs) into immersive VR to create medically coherent virtual patients with distinct, consistent personalities, built on a modular architecture that decouples personality from clinical data. We evaluated our system in a mixed-method, within-subjects study with licensed physicians who engaged in simulated consultations. Results demonstrate that the approach is not only feasible but is also perceived by physicians as a highly rewarding and effective training enhancement. Furthermore, our analysis uncovers critical design principles, including a “realism-verbosity paradox” where less communicative agents can seem more artificial, and the need for challenges to be perceived as authentic to be instructive. This work provides a validated framework and key insights for developing the next generation of socially intelligent VR training environments.
@@ Introduction
Technological advances in immersive environments have transformed virtual reality into a powerful platform capable of reproducing highly realistic physical environments and immersive scenarios. By integrating high-fidelity graphics, motion tracking, and responsive interaction design, virtual reality has found successful applications across domains such as industrial training [37], security [35] and healthcare [9]. Its ability to replicate spatial, visual, and procedural aspects of real-world scenarios with high accuracy has made it particularly valuable in contexts that require experiential learning under safe and controlled conditions, positioning virtual reality as a promising solution for skill acquisition, procedural rehearsal, and decision-making practice where realism is essential to effective transfer of learning [15].
@@
While these recent advances in rendering techniques have markedly improved the visual and physical fidelity of VR systems, the psychological, emotional, and social dimensions of digital humans remain comparatively underexplored [5]. Realistic textures, animations, and gestures can enhance immersion, yet interactions with digital humans are not defined solely by appearance. Human communication inherently involves complex social, emotional, and psychological processes, among which personality plays a fundamental role in shaping how individuals convey information, respond to others, and co-construct meaning during interaction [16, 7]. The absence of mechanisms to represent diverse and consistent personalities risks VR simulations to produce environments that are visually compelling but socially impoverished, thereby limiting their effectiveness in scenarios that rely on authentic interpersonal dynamics.
@@
This gap becomes particularly critical when VR is used to prepare individuals for real-world sociable interactions. In medical consultations, for instance, doctors must not only identify symptoms and apply clinical reasoning but also adapt their communication strategies to patients who differ widely in personality, emotional expressiveness, openness, or levels of trust. Personality shapes how patients present symptoms, respond to questioning, and engage emotionally during clinical encounters [13]. As a result, simulations that neglect these crucial dimensions fail to replicate the full spectrum of a clinical encounter, limiting their effectiveness in developing the nuanced interpersonal competencies essential for effective patient care.
@@
In parallel, recent advancements in natural language process with the advents of large language models (LLMs) has open a promising pathway to simulate human behavior in digital environnments [47]. These models has demonstrated strong capabilities in producing coherent, contextually appropriate, and personality-consistent dialogue [36], allowing move from rigid scripts to dynamic and adaptive communication. Embedding this technology into VR environments has the potential to overcome the current lack of realistic personality in digital humans, unlocking a new level of social faithfulness that is crucial to experiment interpersonal interactions.
@@
Building on prior evidence that patient personality shapes communication patterns in clinical encounters [38, 32, 24], we posit the following hypotheses:
H1: It is feasible to simulate patient with different personality profiles in virtual agents using large language models embedded in immersive VR environments, maintaining both personality consistency and disease symptoms coherence.
@@
(3) Our study provides empirical evidence that personality is a crucial factor in designing effective VR training environments, with significant implications for both medical education and human-computer interaction.
@@ 1.1 Virtual Reality in Medical Training
Virtual reality (VR) has become an established tool in medical education, offering immersive and controlled environments where learners can acquire procedural, diagnostic, and communication skills without risks to real patients [22, 10, 40, 30]. Numerous studies have demonstrated its effectiveness in domains such as surgical training, anatomy learning, and interprofessional collaboration, often reporting improvements in skill acquisition and learner engagement compared to traditional methods [29, 43, 31]. VR-based scenarios have also been used to train delicate competencies, including empathy and breaking bad news, showing that simulated encounters with virtual humans can elicit measurable behavioral changes [14, 4].
In recent years, artificial intelligence (AI) has been integrated into VR training systems to extend their adaptability and realism. AI-driven virtual patients and tutors can generate dynamic feedback, tailor scenarios to individual learner performance, and support assessment of communication quality [2, 20]. For example, LLM-powered agents have been proposed to evaluate clinical decision-making, provide formative feedback, or simulate standardized patients at scale [6, 45]. Despite these advances, most AI-enhanced VR applications remain focused on procedural or technical training (e.g., orthopedics, emergency protocols), with relatively limited exploration of psychosocial and interpersonal challenges, dimensions that are critical in real-world consultations. This gap motivates the need to examine how AI-driven virtual patients can embody nuanced traits such as personality to enrich social realism in VR medical training.
@@ 1.2 Personality Simulation in Digital Agents
Parallel to advances in VR, a growing body of work has investigated how digital agents, particularly those powered by large language models, can be endowed with distinct personalities and communication styles. Personality expression has been shown to significantly influence engagement, trust, and learning outcomes in educational agents [42, 48]. Recent frameworks such as PATIENTSIM [23] and Adaptive-VP [25] highlight the potential of LLMs to simulate clinically relevant personas, ranging from anxious or impatient patients to individuals with limited recall or language proficiency. These systems move beyond scripted dialogues by capturing variation in emotional expression, conversational tone, and interaction difficulty, thereby creating more authentic and challenging training opportunities. Similarly, other works have modeled “difficult” communication styles, such as the Satir-based accuser or rationalizer, to expose learners to demanding interpersonal dynamics [1].
Outside the medical domain, personality-driven agents have also been employed for training in high-stakes contexts such as firefighting and disaster response, where realistic interpersonal stressors and decision-making under pressure are essential [33]. These studies consistently demonstrate that embedding personality traits in digital agents enriches user engagement and promotes transfer of social skills. However, while the impact of personality has been explored in LLM-driven chat-based agents and 2D training systems, its systematic integration into immersive VR medical training remains underdeveloped. This underscores the importance of investigating how patient personality in VR consultations shapes doctors’ strategies, perceptions of realism, and educational outcomes.
Figure 1: The architecture of our interactive patient simulation pipeline. The Virtual Patient Modeling stage (left) uitilizes a real patient inquiry to generate a medically and behaviorally coherent patient. During the Physician Interaction stage (right), a patient response is generated using an LLM that is synthesized into speech via a Text-to-Speech (TTS) module. The corresponding Speech-to-Text (STT) module, which processes the physician’s audio input, is omitted for visual clarity.
@@ 2.1 VR Environment
To provide a realistic and immersive setting for the simulated consultations, we implemented a high-fidelity virtual consultation environment using Unreal Engine, a widely adopted platform for interactive VR applications across domains such as training, simulation, and entertainment [41, 44]. Unreal was selected due to its advanced rendering pipeline, robust support for real-time physics, and seamless integration with head-mounted displays.
The consultation room was modeled after a standard outpatient office, including furniture, medical equipment, and ambient lighting. These design decisions were made to enhance the realism of the simulation and to encourage participants to behave as they would in an actual clinical encounter. Virtual patients were embodied as avatars whose demographic attributes were systematically varied. This strategy aimed to minimize the confounding influence of visual appearance on physicians’ interaction strategies, as prior research can reinforce existing biases and limit learning effectiveness [46].
Naturalistic doctor–patient conversations were enabled through an integrated speech and dialogue pipeline. Doctors’ verbal input was automatically transcribed in real time by a speech-to-text module, after which the transcriptions were processed by a large language model to generate contextually appropriate responses conditioned on both the clinical case and the patient’s personality profile. The responses were then converted into expressive speech using a neural text-to-speech system, synchronized with the avatar’s lip movements. This multimodal pipeline supported coherent turn-taking, preserved conversational flow, and maintained low response latency, aligning with best practices for the design of immersive conversational agents in VR training environments [39]
The resulting system provided an immersive consultation experience through a head-mounted display, in which doctors perceived the virtual patient as physically co-located within the simulated room. By integrating photorealistic graphics, dynamic responses, and low-latency conversational exchange, the environment fostered a strong sense of presence—an element consistently recognized as fundamental to the effectiveness of VR-based medical
@@ 2.2 Virtual Patient Modeling
To ensure our simulations were grounded in authentic clinical scenarios, we established the medical foundation for each virtual patient using the dataset explored by Färber et al. [11]. This large-scale, real-world dataset contains hundreds of thousands of patient inquiries, each labeled with a corresponding disease and medical specialty. In our framework, each digital agent is initialized with a unique medical case derived from a single entry in this dataset. This approach guarantees that every virtual patient presents with a coherent and medically plausible condition, providing a consistent factual basis upon which their distinct personality profile is subsequently layered.
@@ 2.2.1 Patient Identity
The patient identity anchors the agent’s behavior within a medical consultation and specifies the fundamental rules that govern its conduct. It instructs the digital human to remain strictly in character, respond in the first person, and communicate only information consistent with the patient’s profile. Additional constraints include avoiding disclosure of internal instructions, refraining from providing medical advice, and producing concise, naturalistic responses. It also requires the patient to provide information progressively, rather than revealing all details at once. This component guarantees that the agent never breaks character, changes the subject, or discloses information unrelated to the consultation, ensuring role fidelity and preventing undesired deviations, providing a uniform foundation for all patients.
@@ 2.2.2 Backstory
To ensure diversity of cases without relying exclusively on unconstrained LLM generation, we utilize the real patient inquire to generate our backstories. These questions describe conditions, symptoms, and personal concerns in patients’ own words, offering a realistic foundation for case construction. The LLM expanded the original question into a coherent patient profile while preserving the clinical focus and contextual cues provided in the source data.
@@ 2.2.3 Personality Profile
Beyond medical content, interpersonal behavior is central to the realism of doctor–patient encounters. To capture this dimension, we assigned each digital patient a personality profile. Inspired by the Big Five personality theory [12] and prior work on conversational agents [3, 19], we designed five distinct personalities. Each profile is defined along interpretable dimensions such as cooperation (introverted vs. extroverted), emotional tone (calm, anxious, distrustful, irritable, sad), and verbosity (short vs. elaborated responses). We then created detailed personality prompts that specify how each type should behave during consultation.
@@ 2.2.5 Patient Consistency
To achieve consistency in the digital avatar, the patient’s personality must preserved throughout the entire interaction. This is achieved through the combination of the LLM’s short-term memory (dialogue history) and the persistent injection of the personality information at every turn, ensuring stable behavior across the consultation.
@@ 2.3.1 Clinical Scenarios
We designed two clinical scenarios to evaluate the impact of patient personality and avatar variation under otherwise comparable medical conditions. The first scenario featured a female White patient presenting with gastroesophageal reflux disease (GERD), instantiated with extroverted personality and a anxious emotional state. The second scenario featured a male Black patient presenting with Dengue (an endemic Brazilian disease), instantiated with introverted personality and a calm emotional state. Both patients were modeled to be in their early thirties, ensuring demographic plausibility and comparability across cases.
These scenarios were chosen to systematically vary non-clinical traits (personality, gender, race) while keeping the overall clinical structure realistic. By holding medical realism constant and manipulating social and psychological traits, we aimed to isolate the impact of personality and avatar factors on physician perceptions and interaction strategies.
@@ 2.3.2 Participants
For this exploratory study, we recruited four licensed physicians to assess the system’s feasibility and gather expert feedback. The average age was 29.75 years (min = 27, max = 37), and their average clinical experience was 4 years (min = 2, max = 5), and includes specialists in ophthalmology, psychiatry and clinical medical. The sole inclusion criterion was that participants must be actively practicing physicians. The study was conducted in accordance with the approval from our institutional ethics committee, and all participants provided written informed consent prior to their participation.
@@ 2.3.4 Measures
  • Pre-interaction: Before the experiment start, a Simulator Sickness Questionnaire (SSQ) [21] and a baseline familiarity with AI and VR questionary was applied to each participant.
  • Mid-interaction: After each clinical scenario interaction, a questionary was applied to evaluate the dialogue quality, coherence, realism, and perceived personality traits for that patient, totalizing two mid-interaction questionaries per participant.
  • Post-interaction: After the participant interact with both patients, we applied the a SSQ reassessment, User Engagement Scale (UES-SF) [34], NASA-TLX workload assessment [17], VR presence and realism ratings, and a comparative evaluation between the two patients.

### v2

@@ Abstract
While virtual reality (VR) excels at simulating physical environments, its effectiveness for training complex interpersonal skills is limited by a lack of psychologically plausible virtual humans. This gap is particularly critical in medical education, where communication is a core clinical competency. This paper introduces a framework that integrates large language models (LLMs) into immersive VR to create medically coherent virtual patients with distinct, consistent personalities, based on a modular architecture that decouples personality from clinical data. We evaluated the system in a mixed-methods, within-subjects study with licensed physicians conducting simulated consultations. Results suggest that the approach is feasible and perceived as a rewarding and effective training enhancement. Our analysis highlights key design principles, including a “realism–verbosity paradox” and the importance of challenges being perceived as clinically authentic to support learning.
@@ Introduction
Technological advances in immersive environments have transformed virtual reality into a powerful platform capable of reproducing highly realistic physical environments and immersive scenarios. By integrating high-fidelity graphics, motion tracking, and responsive interaction design, virtual reality has found successful applications across domains such as industrial training [radhakrishnan2021systematic], security [10257288] and healthcare [ekstrand2018immersivem]. Its ability to replicate spatial, visual, and procedural aspects of real-world scenarios with high accuracy has made it particularly valuable in contexts that require experiential learning under safe and controlled conditions, positioning virtual reality as a promising solution for skill acquisition, procedural rehearsal, and decision-making practice where realism is essential to effective transfer of learning [gunn2018use].
@@
While these recent advances in rendering techniques have markedly improved the visual and physical fidelity of VR systems, the psychological, emotional, and social dimensions of digital humans remain comparatively underexplored [vcekic2025virtual]. Realistic textures, animations, and gestures can enhance immersion, yet interactions with digital humans are not defined solely by appearance. Human communication inherently involves complex social, emotional, and psychological processes, among which personality plays a fundamental role in shaping how individuals convey information, respond to others, and co-construct meaning during interaction [hassan2019anatomy, dhillon2023impact]. The absence of mechanisms to represent diverse and consistent personalities risks VR simulations to producing environments that are visually compelling but socially impoverished, thereby limiting their effectiveness in scenarios that rely on authentic interpersonal dynamics.
@@
This gap becomes particularly critical when VR is used to prepare individuals for real-world sociable interactions. In medical consultations, for instance, doctors must not only identify symptoms and apply clinical reasoning but also adapt their communication strategies to patients who differ widely in personality, emotional expressiveness, openness, or levels of trust. Personality shapes how patients present symptoms, respond to questioning, and engage emotionally during clinical encounters [flynn2007personality]. As a result, simulations that neglect these crucial dimensions fail to replicate the full spectrum of a clinical encounter, limiting their effectiveness in developing the nuanced interpersonal competencies essential for effective patient care.
@@
In parallel, recent advancements in natural language processing with the advent of large language models (LLMs) have opened a promising pathway to simulate human behavior in digital environments [xie2024can]. These models have demonstrated strong capabilities in producing coherent, contextually appropriate, and personality-consistent dialogue [park2023generative], allowing a move from rigid scripts to dynamic and adaptive communication. Embedding this technology into VR environments has the potential to overcome the current lack of realistic personality in digital humans, unlocking a new level of social faithfulness that is crucial for studying interpersonal interactions.
@@
Building on prior evidence that patient personality shapes communication patterns in clinical encounters [redelmeier2021understanding, oldham2020personality, lee2025personality], we posit the following hypotheses:
H1: It is feasible to simulate patients with different personality profiles in virtual agents using large language models embedded in immersive VR environments, maintaining both behavioral consistency and coherence of disease symptoms.
@@
(3) Our pilot study provides preliminary evidence suggesting that personality is a significant factor in designing effective VR training environments, with significant implications for both medical education and human-computer interaction.
@@ 1.1 Virtual Reality in Medical Training
Virtual reality (VR) has become an established tool in medical education, offering immersive and controlled environments where learners can acquire procedural, diagnostic, and communication skills without risks to real patients [kyaw2019virtual, ekstrand2018immersivem, sattar2019effects, muhling2025personality]. Numerous studies have demonstrated its effectiveness in domains such as surgical training, anatomy learning, and interprofessional collaboration, often reporting improvements in skill acquisition and learner engagement compared to traditional methods [mergen2024reviewing, sung2024effectiveness, muhling2025comparing]. VR-based scenarios have also been used to train delicate competencies, including empathy and breaking bad news, showing that simulated encounters with virtual humans can elicit measurable behavioral changes [guetterman2019medical, carnell2022informing].
In recent years, artificial intelligence (AI) has been integrated into VR training systems to extend their adaptability and realism. AI-driven virtual patients and tutors can generate dynamic feedback, tailor scenarios to individual learner performance, and support assessment of communication quality [borg2025virtual, kapadia2024evaluation]. For example, LLM-powered agents have been proposed to evaluate clinical decision-making, provide formative feedback, or simulate standardized patients at scale [cook2025virtual, weisman2025development]. Despite these advances, most AI-enhanced VR applications remain focused on procedural or technical training (e.g., orthopedics, emergency protocols), with relatively limited exploration of psychosocial and interpersonal challenges, dimensions that are critical in real-world consultations. This gap motivates the need to examine how AI-driven virtual patients can embody nuanced traits such as personality to enrich social realism in VR medical training.
@@ 1.2 Personality Simulation in Digital Agents
Parallel to advances in VR, a growing body of work has investigated how digital agents, particularly those powered by large language models, can be endowed with distinct personalities and communication styles. Personality expression has been shown to significantly influence engagement, trust, and learning outcomes in educational agents [sonlu2024effects, zhu2025social]. Recent frameworks such as PATIENTSIM [kyung2025patientsim] and Adaptive-VP [lee2025adaptive] highlight the potential of LLMs to simulate clinically relevant personas, ranging from anxious or impatient patients to individuals with limited recall or language proficiency. These systems move beyond scripted dialogues by capturing variation in emotional expression, conversational tone, and interaction difficulty, thereby creating more authentic and challenging training opportunities. Similarly, other works have modeled “difficult” communication styles, such as the Satir-based accuser or rationalizer, to expose learners to demanding interpersonal dynamics [bodonhelyi2025modeling].
Outside the medical domain, personality-driven agents have also been employed for training in high-stakes contexts such as firefighting and disaster response, where realistic interpersonal stressors and decision-making under pressure are essential [oliveira2024exploring]. These studies consistently demonstrate that embedding personality traits in digital agents enriches user engagement and promotes transfer of social skills. However, while the impact of personality has been explored in LLM-driven chat-based agents and 2D training systems, its systematic integration into immersive VR medical training remains underdeveloped. This underscores the importance of investigating how patient personality in VR consultations shapes doctors’ strategies, perceptions of realism, and educational outcomes.
Figure 1: The architecture of our interactive patient simulation pipeline. The Virtual Patient Modeling stage (left) utilizes a real patient inquiry to generate a medically and behaviorally coherent patient. During the Physician Interaction stage (right), a patient response is generated using an LLM that is synthesized into speech via a Text-to-Speech (TTS) module. The corresponding Speech-to-Text (STT) module, which processes the physician’s audio input, is omitted for visual clarity.
@@ 2.1 VR Environment
To provide a realistic and immersive setting for the simulated consultations, we implemented a high-fidelity virtual consultation environment using Unreal Engine, a widely adopted platform for interactive VR applications across domains such as training, simulation, and entertainment [schmitt2022nureality, vercelli2024risk]. Unreal was selected due to its advanced rendering pipeline, robust support for real-time physics, and seamless integration with head-mounted displays.
The consultation room was modeled after a standard outpatient office, including furniture, medical equipment, and ambient lighting. These design decisions were made to enhance the realism of the simulation and to encourage participants to behave as they would in an actual clinical encounter. Virtual patients were embodied as avatars whose demographic attributes were systematically varied. This strategy aimed to minimize the confounding influence of visual appearance on physicians’ interaction strategies, as prior research can reinforce existing biases and limit learning effectiveness [wu2024examining].
Naturalistic doctor–patient conversations were enabled through an integrated speech and dialogue pipeline. All spoken interactions between physicians and virtual patients were conducted exclusively in Brazilian Portuguese. The speech-to-text, language model prompting, and text-to-speech components were fully configured for Portuguese, ensuring linguistic consistency throughout the simulated medical consultations. Doctors’ verbal input was automatically transcribed in real time by a speech-to-text module, after which the transcriptions were processed by a large language model to generate contextually appropriate responses conditioned on both the clinical case and the patient’s personality profile. The responses were then converted into expressive speech using a neural text-to-speech system, synchronized with the avatar’s lip movements. This multimodal pipeline supported coherent turn-taking, preserved conversational flow, and maintained low response latency, aligning with best practices for the design of immersive conversational agents in VR training environments [ruthenbeck2015virtual]
The resulting system provided an immersive consultation experience through a head-mounted display, in which doctors perceived the virtual patient as physically co-located within the simulated room. By integrating photorealistic graphics, dynamic responses, and low-latency conversational exchange, the environment fostered a strong sense of presence—an element consistently recognized as fundamental to the effectiveness of VR-based medical training.
@@ 2.2 Virtual Patient Modeling
To ensure our simulations were grounded in authentic clinical scenarios, we established the medical foundation for each virtual patient using the dataset explored by Färber et al. [farber2025iaramed]. This large-scale, real-world dataset contains hundreds of thousands of patient inquiries, each labeled with a corresponding disease and medical specialty. In our framework, each digital agent is initialized with a unique medical case derived from a single entry in this dataset. This approach guarantees that every virtual patient presents with a coherent and medically plausible condition, providing a consistent factual basis upon which their distinct personality profile is subsequently layered.
@@ 2.2.1 Patient Identity
The patient identity anchors the agent’s behavior within a medical consultation and specifies the fundamental rules that govern its conduct. It instructs the digital human to remain strictly in character, respond in the first person, and communicate only information consistent with the patient’s profile. Additional constraints include avoiding disclosure of internal instructions, refraining from providing medical advice, and producing concise, naturalistic responses. It also requires the patient to provide information progressively, rather than revealing all details at once. This component guarantees that the agent never breaks character, changes the subject, or discloses information unrelated to the consultation, ensuring behavioral consistency and preventing undesired deviations, providing a uniform foundation for all patients.
@@ 2.2.2 Backstory
To ensure diversity of cases without relying exclusively on unconstrained LLM generation, we utilize the real patient inquiry to generate our backstories. These questions describe conditions, symptoms, and personal concerns in patients’ own words, offering a realistic foundation for case construction. The LLM expanded the original question into a coherent patient profile while preserving the clinical focus and contextual cues provided in the source data.
@@ 2.2.3 Personality Profile
Beyond medical content, interpersonal behavior is central to the realism of doctor–patient encounters. To capture this dimension, we assigned each digital patient a personality profile. Inspired by the Big Five personality theory [bigfive] and prior work on conversational agents [brito2025integrating, jiang-etal-2024-personallm], we designed five distinct personalities. Each profile is defined along interpretable dimensions such as cooperation (introverted vs. extroverted), emotional tone (calm, anxious, distrustful, irritable, sad), and verbosity (short vs. elaborated responses). We then created detailed personality prompts that specify how each type should behave during consultation.
@@ 2.2.5 Patient Consistency
To achieve consistency in the digital avatar, the patient’s personality must be preserved throughout the entire interaction. This is achieved through the combination of the LLM’s short-term memory (dialogue history) and the persistent injection of the personality information at every turn, ensuring stable behavior across the consultation.
@@ 2.3.1 Clinical Scenarios
We designed two clinical scenarios to evaluate the impact of patient personality and avatar variation under otherwise comparable medical conditions. The first scenario featured a female White patient presenting with gastroesophageal reflux disease (GERD), instantiated with extroverted personality and an anxious emotional state. The second scenario featured a male Black patient presenting with Dengue (an endemic Brazilian disease), instantiated with introverted personality and a calm emotional state. Both patients were modeled to be in their early thirties, ensuring demographic plausibility and comparability across cases.
These scenarios were designed to explore the combined influence of personality and avatar variation under realistic clinical conditions. By holding medical realism constant and manipulating social and psychological traits, we aimed to explore the combined impact of personality and avatar factors on physician perceptions and interaction strategies.
@@ 2.3.2 Participants
For this exploratory study, we recruited four licensed physicians to assess the system’s feasibility and gather expert feedback. The average age was 29.75 years (min = 27, max = 37), and their average clinical experience was 4 years (min = 2, max = 5), and include specialists in ophthalmology, psychiatry and general medicine. The sole inclusion criterion was that participants must be actively practicing physicians. The study was conducted in accordance with approval from our institutional ethics committee, and all participants provided written informed consent prior to their participation.
@@ 2.3.4 Measures
  • Pre-interaction: Before the experiment start, the Simulator Sickness Questionnaire (SSQ) [kennedy1993simulator] and a baseline familiarity with AI and VR questionnaire was applied to each participant.
  • Mid-interaction: After each clinical scenario interaction, a questionnaire was applied to evaluate the dialogue quality, coherence, realism, and perceived personality traits for that patient, resulting in two mid-interaction questionnaires per participant.
  • Post-interaction: After the participant interacted with both patients, we applied an SSQ reassessment, User Engagement Scale (UES-SF) [o2018practical], NASA-TLX workload assessment [hoonakker2011measuring], VR presence and realism ratings, and a comparative evaluation between the two patients.